OpenStudy (anonymous):

Write an equation for a circle with center at (2,3) and radius of 6

OpenStudy (aakashsudhakar):

The standard equation for a circle is... \[(x - h)^2 + (y - k)^2 = r^2\] You're given that (h, k) = (2, 3) and that r = 6. Simply plug these values in and solve for the proper equation of the circle. (x - h)^2 + (y - k)^2 = r^2 (h, k) --> (2, 3) r = 6 (x - 2)^2 + (y - 3)^2 = 36 Therefore, your circle equation is... \[(x - 2)^2 + (y - 3)^2 = 36\]

jigglypuff314 (jigglypuff314):

@AakashSudhakar it is against this site's code of conduct to give answers, so please try to keep that in mind from now on :) http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

its fine because the person showed me how to solve it because ive been working on this once before

OpenStudy (aakashsudhakar):

@jigglypuff314: My apologies; I'll refrain from disclosing answers from this point on.